OpenStudy (anonymous):

If a = 4, then a 2 • a 3 is equivalent to all of the following except... a) 4^6 b) a^5 c) 1,024 d) 4^2 • 4^3

OpenStudy (anonymous):

*a^2 • a^3

OpenStudy (anonymous):

a^2 • a^3 = 4^2 • 4^3. We can rewrite that as 4^2+3 = 4^5 4^5 is also equal to 1024. So.... A is the only answer that doesn't make sense.